OpenStudy (anonymous):

help:) f(x) = 3x-1 g(x)= 5/2x+7 h(x)= x^2+1 find:(fg/h)(-1)

OpenStudy (anonymous):

are you sure that this question is correct?

OpenStudy (anonymous):

the rquired is f(-1)g(-1)/h(-1)=(-3-1)(-5/2+7)/(1+1)=-4*9/2*2=-9 hope its correct

OpenStudy (anonymous):

yah.. mehmetzeki